What are the locations of the vertices: (x-4)^2/25 - (y-7)^2/9 = 1
Answer by stanbon(75887)

What are the locations of the vertices: (x-4)^2/25 - (y-7)^2/9 = 1
--------
Center at (4,7)
a = sqrt(25) = 5
----
Vertices:
(4-5,7) = (-1,7)
(4+5,7) = (9,7)
===================
cheers,
Stan H.